johnny is 24 years old, which is three years less than 3 times joeys age.
how old is joey.

Answer by Maths68(1474)

Let
Johnny Present age = j
Joeys Present age = k
Johnny is 24 years old means we have to add 24 to the present age of Joeys to equalize their ages
j=k+24............(1)
which is three years less than 3 times Joeys age
Johnny Present age = 3 times Joeys age - 3
j=3k-3............(2)
Put the value of j from (1) to (2)
k+24=3k-3
k-3k=-3-24
-2k=-27
-2k/-2=-27/2
k=13.5
Put the value of k in (1)
j=k+24
j=13.5+24
j=37.5

Johnny Present age = j = 37 and half years old
Joeys Present age = k = 13 and half years old